# ContextQFormer: A New Context Modeling Method for Multi-Turn Multi-Modal Conversations

### v1

@@ 2.3.   Multi-Modal Dialogue Dataset
Dialogues have received continuous attention in the field of NLP for many years, and with the development of multi-modal, image-based multi-modal dialogues have gradually gained attention, such as VisualDialogueDas et al. (2017), PhotoChatZang et al. (2021) and MMDialogueFeng et al. (2022). VisualDialogue mainly solves multiple rounds of Q&A tasks, while PhotoChat focuses more on whether to share an image in the conversation scene. MMDialogue collects images and conversations from social media. However, these data are not built for multi-modal large language models, so the dependency of multi-modal large models on data cannot be met in terms of dialogue rounds and task complexity.
@@ 3.1.   Training Dataset
TMDialog-IT. TMDialog-IT is utilized for instruction tuning, which includes three components with a total amount of data exceeding 1.5 million. The first part consists of publicly available multi-modal multi-turn conversation datasets including MiniGPT-4 Zhu et al. (2023), LLAVA Liu et al. (2023), VQA-v2 Goyal et al. (2017) and some unpublished data. The second part consists of a natural language multi-turn conversation dataset, including MOSS Sun et al. (2023). Lastly, we generated our own dataset utilizing GPT-4 API including three categories: Interaction, Long Memory, and Multi Images. The precise definitions of these types of data are provided in Table 1. The dataset generation process is shown in figure 2.
@@ 8.   Bibliographical References
  • Liu et al. (2023) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2023. Visual instruction tuning. arXiv preprint arXiv:2304.08485.

### v2

@@ 2.3.   Multi-Modal Dialogue Dataset
Dialogues have received continuous attention in the field of NLP for many years, due to their fundamental role in natural language understanding and generation. In recent years, several high-quality textual dialogue datasets have been developed to support multilingual Liu et al. (2023b), recommendation-based Liu et al. (2020), and medical Shi et al. (2024) dialogue scenarios. While these datasets have expanded the scope of dialogue research, they are still restricted to text-only interactions. With the development of multimodal learning, image-based multimodal dialogues have gradually gained attention, such as VisualDialogue Das et al. (2017), PhotoChat Zang et al. (2021), and MMDialogue Feng et al. (2022). VisualDialogue mainly solves multiple rounds of Q&A tasks, while PhotoChat focuses more on whether to share an image in the conversation scene. MMDialogue collects images and conversations from social media. However, these data are not built for multimodal large language models, so the dependency of multimodal large models on data cannot be met in terms of dialogue rounds and task complexity.
@@ 3.1.   Training Dataset
TMDialog-IT. TMDialog-IT is utilized for instruction tuning, which includes three components with a total amount of data exceeding 1.5 million. The first part consists of publicly available multi-modal multi-turn conversation datasets including MiniGPT-4 Zhu et al. (2023), LLAVA Liu et al. (2023a), VQA-v2 Goyal et al. (2017) and some unpublished data. The second part consists of a natural language multi-turn conversation dataset, including MOSS Sun et al. (2023). Lastly, we generated our own dataset utilizing GPT-4 API including three categories: Interaction, Long Memory, and Multi Images. The precise definitions of these types of data are provided in Table 1. The dataset generation process is shown in figure 2.
@@ 8.   Bibliographical References
  • Liu et al. (2023a) Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. 2023a. Visual instruction tuning. arXiv preprint arXiv:2304.08485.
@@
  • Liu et al. (2023b) Zeming Liu, Ping Nie, Jie Cai, Haifeng Wang, Zheng-Yu Niu, Peng Zhang, Mrinmaya Sachan, and Kaiping Peng. 2023b. XDailyDialog: A multilingual parallel dialogue corpus. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 12240–12253, Toronto, Canada. Association for Computational Linguistics.
  • Liu et al. (2020) Zeming Liu, Haifeng Wang, Zheng-Yu Niu, Hua Wu, Wanxiang Che, and Ting Liu. 2020. Towards conversational recommendation over multi-type dialogs. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 1036–1049, Online. Association for Computational Linguistics.
@@
  • Shi et al. (2024) Xiaoming Shi, Zeming Liu, Li Du, Yuxuan Wang, Hongru Wang, Yuhang Guo, Tong Ruan, Jie Xu, Xiaofan Zhang, and Shaoting Zhang. 2024. Medical dialogue system: A survey of categories, methods, evaluation and challenges. In Findings of the Association for Computational Linguistics: ACL 2024, pages 2840–2861, Bangkok, Thailand. Association for Computational Linguistics.